we have a pile of dirt with the measurements of 26 ft wide base, 100 ft long, and 8 ft high, how do we measure this and turn it into cubic yards, this is more of a right triangle than triangle
Answer by mananth(16946)

we have a pile of dirt with the measurements of 26 ft wide base, 100 ft long, and 8 ft high, how do we measure this and turn it into cubic yards, this is more of a right triangle than triangle
3 feett = 1yard
26 feet = 26/3 = 8.66 yards
..
100 feet = 100/3= 33.33 yards
..
8 feet = 8 / 3 = 2.66 yards
..
Volume = 8.66*33.33*2.66
= 767.77 cu. yards
..
This cannot be a triangle . triangle is two dimensional.
this can be represented as a cuboid